OpenStudy (anonymous):

A two-way frequency table shows grades for students in college and students in high school: High school College Total GPA above 3.0 14 26 40 GPA below 3.0 46 14 60 60 40 100 Based on this data, are "being in high school" and "GPA above 3.0" independent events? Yes, P(high school | GPA above 3.0) = P(high school) Yes, P(high school | GPA above 3.0) = P(GPA above 3.0) No, P(high school | GPA above 3.0) ≠ P(high school) No, P(high school | GPA above 3.0) ≠ P(GPA above 3.0) @satellite73

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hello friend :)

OpenStudy (anonymous):

let me get on a real computer so i can look more carefully at the question

OpenStudy (anonymous):

hahah okay

OpenStudy (anonymous):

\[p(high~school)=\frac{14}{100}=.14\]right?

OpenStudy (anonymous):

which one? @satellite73

OpenStudy (anonymous):

oh no that is not right, it is hard for me to read the table the total number of people in high school is 60 the way i read it

OpenStudy (anonymous):

and the total all together is 100 to the probability of high school is \(\frac{60}{100}=0.6\)

OpenStudy (anonymous):

High school College Total GPA above 3.0 14 26 40 GPA below 3.0 46 14 60 60 40 100

OpenStudy (anonymous):

then the number of people with GPA over 3.0 is 40 of those, 14 are in high school so \[p(high~school|gpa>3.0)=\frac{14}{40}\]

OpenStudy (anonymous):

so the answer is B?

OpenStudy (anonymous):

or A?

OpenStudy (anonymous):

and since \[\frac{14}{40}\neq \frac{60}{100}\] the events are Depenedent

OpenStudy (anonymous):

they are not the same, so C

OpenStudy (anonymous):

oh thx

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

oops i mean "you're welcome"

OpenStudy (anonymous):

hahaha it's fine.

OpenStudy (anonymous):

A political candidate wanted to understand what issues were important to their voters. They asked 1,999 voters, "Which issue is the most important: crime, city planning, or health care?" The results of the survey are shown in the two-way table below: Crime City Planning Health Care Male 118 239 458 Female 418 317 449 What is the probability that a person chosen at random from this survey is concerned about crime, given that they are male? Round your answer to the nearest tenth. 5.9% 14.5% 22.0% 26.8% @satellite73

OpenStudy (anonymous):

total male is 458 right?

OpenStudy (anonymous):

of those 118 concerned about crime, so \[118\div 458\]

OpenStudy (anonymous):

.26

OpenStudy (anonymous):

wait i dont get any of those answers

OpenStudy (anonymous):

crap maybe i am reading the table wrong i get like 25.7

OpenStudy (anonymous):

well i don't know what else to do

OpenStudy (anonymous):

so it would be D cuz the question asks "round your answer to the nearest tenth"

OpenStudy (anonymous):

should be 25.8 in that case

OpenStudy (anonymous):

not26.8

OpenStudy (anonymous):

but D is the closest

OpenStudy (anonymous):

oh let me send u a good table

OpenStudy (anonymous):

screen shot?

OpenStudy (anonymous):

Crime City Planning Health Care Male 118 239 458 Female 418 317 449 @satellite73

OpenStudy (anonymous):

yeah i thought so, still 118/458

OpenStudy (anonymous):

pick D i guess

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got it wrong i got a 60% on the assignment

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i am sticking with my answer on that one do you have to do it again?

OpenStudy (anonymous):

its up to me but i wanna do it again cuz im not happy with the grade.

OpenStudy (anonymous):

you are a worker

OpenStudy (anonymous):

hahah thx i guess.

OpenStudy (anonymous):

A compact disc is designed to last an average of 4 years with a standard deviation of 0.9 years. What is the probability that a CD will last less than 3 years? 1.11% 13.35% 86.65% 100% @satellite73

OpenStudy (anonymous):

oh my

OpenStudy (anonymous):

you need some kind of calculator or table for this, you can't do this in your head but it is probably the second one, but maybe the first let me see if i can google a normal curve calculator

OpenStudy (anonymous):

so u cant do this?

OpenStudy (anonymous):

sure i can, but i need a table of some kind or a calculator give me a second i think i have one

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

http://stattrek.com/online-calculator/normal.aspx

OpenStudy (anonymous):

is 13.3% an answer

OpenStudy (anonymous):

yeah it is B like i said, you need a calculator

OpenStudy (anonymous):

okay thank you sir i appreciate it sir god bless ya sir i need to stop something is wrong with me okay bye sir LOL

OpenStudy (anonymous):

more sir's than i have had in all my life good luck with the rest of it

OpenStudy (anonymous):

hahahha no sir i still have more i think so dont escape sir.

OpenStudy (anonymous):

ok i will be here watching the man from uncle on tv holla if you need help

OpenStudy (anonymous):

i will hollaaa!! xD